OpenStudy (anonymous):

A piece of paper is 0.1 mm thick. When folded, the paper is twice as thick. Assume that you could fold the paper as many times as you want. How many folds would be required for the paper to be taller than Mount Everest(8850m)?

OpenStudy (anonymous):

solve \[0.1\times 2^x=8850\]

OpenStudy (anonymous):

oh that is wrong sorry

OpenStudy (anonymous):

what is a mm? one thousands of a meter?

OpenStudy (anonymous):

1m=100cm, 1cm=10mm, so...ya.

OpenStudy (anonymous):

so solve \[0.0001\times 2^x=8850\] first step is to divide by \(0.0001\) aka multiply by \(10,000\) and get \(2^x=88,500,000\)

OpenStudy (anonymous):

then solve for x either by trial and error or using logs. do you know how to use logs for this?

OpenStudy (anonymous):

the log method is to write \[x=\frac{\log(88,500,000)}{\log(2)}\]

OpenStudy (anonymous):

why would you multiply by .0001? I thought it was just .1 mm, and that you only convert the 8850 to 8850(000). And no, I don't know how to use logs for this...

OpenStudy (anonymous):

i get around 27, perhaps a surprisingly small number

OpenStudy (anonymous):

ok lets got back to the mm the height of mt everest is given in meters, and the width of the paper is given in mm so we have to convert them to the same units otherwise we cannot compute anything

OpenStudy (anonymous):

how??? I think I get how you went to the equation you did, but how to go from there?

OpenStudy (anonymous):

apparently 1 mm = .001 meters (one thousandth of a meter) and so 0.1 mm = .0001 meters right?

OpenStudy (anonymous):

therefore first we want to convert mm to meters then we know the number of pieces of paper, when folded in half go like this

OpenStudy (anonymous):

\(2, 4, 8, 16, 32, ...\) i.e.. \(2, 2^2,2^3,2^4,...\)

OpenStudy (anonymous):

Oh, yeah!!! I was just doing it the other way. converting to mm instead of m. I just realized that...srry

OpenStudy (anonymous):

so after "n" folds the number of pieces of paper will be \(2^n\)

OpenStudy (anonymous):

each with width .0001 mm so we want to solve \(.0001\times 2^n=8850\) for n

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

Well, I don't get the part when you are showing the 2, 4, and all the other numbers... where did you get the 2? I s that from doubling? And where did the 0.1 go?

OpenStudy (anonymous):

site is crashing on me, so i may not be able to respond, but the idea now is to multiply both sides of the quation by 10000, then see if you can find an \(n\) for which \(2^n\geq 88,500,000\)

OpenStudy (anonymous):

yes you are doubling

OpenStudy (anonymous):

wait, are you using 0.1 with 88500000?

OpenStudy (anonymous):

each time you fold the paper in half you double the number of pieces of paper. as for the .0001 it is still there, i just wrote the numbers \(2, 4, 8, 16, 32...\) so you could see that the paper is doubling and after n folds you will have \(2^n\) layers, each of which is \(.0001 m\) thick, for a total thikness of \(.0001\times 2^n\)

OpenStudy (anonymous):

i got rid of the .1 by conveting it to meters, so now the .1 mm is .0001 m

OpenStudy (anonymous):

that is why i wrote \[.0001\times 2^x=8850\]

OpenStudy (anonymous):

ya, I get that (btw, my computer is not exactly the fastest computer on the face of the earth, so if I ask you a question right after you answer it, it is because of the buffering)

OpenStudy (anonymous):

So I don't get how to solve the log part...

OpenStudy (anonymous):

if you don't knowabout logs yet, then solve by trial and error. see what power you raise 2 to in order to get a number bigger than 88,500,000

OpenStudy (anonymous):

oh, I see... So I tried out the logarithm thing on my calculator and I got about 26.402, and I would round that up, right? And so that would give me 27? But how... I mean, can... you do that logarithm equation by hand?